OpenStudy (anonymous):

Is (3, 13) a solution of the equation y = 4x?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

think of (3,13) as (x,y)

OpenStudy (anonymous):

3,13

OpenStudy (unklerhaukus):

is (13)=4(3) ?

OpenStudy (anonymous):

12(13)

OpenStudy (unklerhaukus):

is (13)=4(3) ? is 13=4×3 ?